Andre rents beach chairs and umbrellas to tourists visiting the beach. On one Saturday, he has 150 customers. If he rents 125 chairs and 72 people rent both chairs and umbrellas, how many people rent umbrellas?
1.)25
2.)72
3.)97
4.)125

Answer by ewatrrr(24785)


Hi
125 + 72 - 150 = 47
125 - 47 = 78 just chairs
72 - 47 = 25 just umbrellas
47 Both
150